# Whirling injections, surjections, and other functions between finite sets

### v4

@@ 1. Introduction
The whirl maps $w_{i}$ are a generalization of the toggle maps that are heavily studied in dynamical algebraic combinatorics. Toggles were first introduced as a way of writing the rowmotion map as a composition of simpler operations on order ideals [CF95]. Striker later generalized them to a much wider variety of settings [Str18]. There have been many interesting homomesy and periodicity phenomena discovered for actions defined in terms of toggles; see e.g., [PR15, Rob16, Had21, EFG+16, JR18, ELM+24].
@@
In Section 2 we consider the whirling action on injections and on surjections from a finite set $[n]$ to another finite set $[k]$. In both cases, we prove that within each orbit, any two elements $1,2,\dots,k$ of the codomain appear as outputs of functions the same number of times. We also prove this for a family of functions that naturally generalizes injections and conjecture it for a similar family that generalizes surjections. Then in Section 3, we explore the whirling action on various families of restricted growth words, which are functions that encode set partitions.
@@ 2. Whirling on m-injections and m-surjections between finite sets
In this section we first define the discrete functions $m$-injections and $m$-surjections and the whirling action on them, along with the basic tools (e.g., orbit boards) that we use to understand the dynamics. The next few sections detail the lemmas and proofs of homomesy. Then we discuss a consequence of the homomesy for orbit sizes of whirling, for which we have no other proof. We finish with a lifting of the dynamics to the piecewise-linear setting, where the homomesy conjecturally appears to continue to hold.
@@ 2.1. The whirling map on m-injections and m-surjections
For $f\in\mathcal{F}\subseteq[k]^{[n]}$ define a map $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$, called whirling at index $i$, as follows: repeatedly add 1 (mod $k$) to the value of $f(i)$ until we get a function in $\mathcal{F}$. The new function is $w_{i}(f)$.
@@
We now discuss orbits and homomesy for $\mathbf{w}$. Theorem 2.10 is the main result of this section.
@@
Conjecture 2.11 has been verified by a computer for all triples $(m,n,k)$ where $n\leq 10$ (for which there are finitely many relevant triples since $\operatorname{Sur}_{m}(n,k)=\O$ if $mk>n$).
@@ 2.2. Orbit board terminology and notation
To explain why the claim is true, we consider two cases
@@
Suppose that we get stuck in partitioning the orbit board into $[k]$-chunks. In this case, when creating a chunk $C$ through $\{1,\dots,i\}$, we have $i$ in position $P$ and all $(i+1)$-entries in $(P,[1,n])$ are already part of chunks. Then we will choose a chunk $C^{\prime}$ for which both the $i$ and $(i+1)$-entries are within $(P,[1,n])$; such a chunk exists by the argument above. We will take the positions of the $i+1,\dots,k$-entries of $C^{\prime}$ and make them part of the chunk $C$ instead of $C^{\prime}$. Then it is just as if $C$ was created previously, and we now must complete $C^{\prime}$. This can cause a chain reaction of chunk reassignment and we will now discuss why the process must end.
@@ 1 $\normalsize{{\color{black}4}}$⃝ 3 $\normalsize{{\color{black}3}}$⃝ 1 2
Any $n$ consecutive positions (P,[0,n1])(P,[0,n-1]) contains the multiset of values of some function in $\operatorname{Sur}_{1}(n,k)$ and thus (P,[1,n1])(P,[1,n-1]) contains a position with entry $i+1$. To prove (1), Lemma 2.19 assures us that it suffices to prove that (Q,[1,n1])(Q,[1,n-1]) does not contain $i+1$ in any position. Every position in (Q,[1,n1])(Q,[1,n-1]) is either
@@
Since $Q$ is the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$, any position simultaneously in both (Q,[1,n1])(Q,[1,n-1]) and (P,[1,n1])(P,[1,n-1]) cannot contain $i+1$. The position below $P$ contains $i$ since $P$ is a stoplight. Let $R$ be a position between $P$ and $Q$. If the position below $R$ contains $i+1$, then $R$ contains either $i$ or $i+1$. Since $P$ is a stoplight, no position in (P,[1,n1])(P,[1,n-1]) contains $i$, so $R$ cannot contain $i$ (which proves (2)). If $R$ contains $i+1$, then $R$ is not a stoplight because $Q$ is a position in (R,[1,n1])(R,[1,n-1]) with the same value as $R$. So the position below $R$ cannot contain $i+1$, meaning no positions in (Q,[1,n1])(Q,[1,n-1]) contain $i+1$.
@@ 3.2. Proof of homomesy for whirling restricted growth words
[…]
@@ 3.3. Whirling on noncrossing RG-words
Let $f\in\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. If f(n1)=f(n)f(n-1)=f(n), then $(w_{n}(f))(i)=1$.

### v5

@@ 1. Introduction
The whirling maps $w_{i}$ are a generalization of the toggle maps that are heavily studied in dynamical algebraic combinatorics. Toggles were first introduced as a way of writing the rowmotion map as a composition of simpler operations on order ideals [CF95]. Striker later generalized them to a much wider variety of settings [Str18]. There have been many interesting homomesy and periodicity phenomena discovered for actions defined in terms of toggles; see e.g., [PR15, Rob16, Had21, EFG+16, JR18, ELM+24].
@@
In Section 2 we consider the whirling action on injections and on surjections from a finite set $[n]$ to another finite set $[k]$. In both cases, we prove that within each orbit $\mathcal{O}$, the number $\sum\limits_{f\in\mathcal{O}}\#f^{-1}(i)$ of times $i\in[k]$ appears as an output of a function in $\mathcal{O}$ is the same for each $i\in[k]$. We also prove this for a family of functions that naturally generalizes injections and conjecture it for a similar family that generalizes surjections. Then in Section 3, we explore the whirling action on various families of restricted growth words, which are functions that encode set partitions.
@@ 2. Whirling on m-injections and m-surjections between finite sets
In this section we first define the families of functions, the $m$-injections and $m$-surjections, that we will consider, and the whirling action on them. We also define the basic tools (e.g., orbit boards) that we use to understand the dynamics. The next few subsections detail the lemmas and proofs of homomesy. Then we discuss a consequence of the homomesy for orbit sizes of whirling, for which we have no other proof. We finish with a lifting of the dynamics to the piecewise-linear setting, where the homomesy conjecturally appears to continue to hold.
@@ 2.1. The whirling map on m-injections and m-surjections
For $f\in\mathcal{F}\subseteq[k]^{[n]}$ we define a map $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$, called whirling at index $i$, as follows: repeatedly add 1 (mod $k$) to the value of $f(i)$ until we get a function in $\mathcal{F}$. The new function is $w_{i}(f)$.
@@
We now discuss orbits and homomesy for $\mathbf{w}$. Theorem 2.10 is the main result of this section. We will prove the different cases of this theorem in the upcoming subsections 2.32.4, and 2.5.
@@
Conjecture 2.11 has been verified by a computer for all triples $(m,n,k)$ where $n\leq 10$ (for which there are finitely many relevant triples since $\operatorname{Sur}_{m}(n,k)=\varnothing$ if $mk>n$).
@@ 2.2. Orbit board terminology and notation
To explain why the claim is true, we consider two cases.
@@
Suppose that we get stuck in partitioning the orbit board into $[k]$-chunks. In this case, when creating a chunk $C$ through $\{1,\dots,i\}$, we have $i$ in position $P$ and all $(i+1)$-entries in $(P,[1,n])$ are already part of chunks. Then we will choose a chunk $C^{\prime}$ for which both the $i$ and $(i+1)$-entries are within $(P,[1,n])$; such a chunk exists by the claim. We will take the positions of the $i+1,\dots,k$-entries of $C^{\prime}$ and make them part of the chunk $C$ instead of $C^{\prime}$. Then it is just as if $C$ was created previously, and we now must complete $C^{\prime}$. This can cause a chain reaction of chunk reassignment and we will now discuss why the process must end.
@@ 1 $\normalsize{{\color{black}4}}$⃝ 3 $\normalsize{{\color{black}3}}$⃝ 1 2
Any set of $n$ consecutive positions (P,[0,n1])(P,[0,n-1]) contains the multiset of values of some function in $\operatorname{Sur}_{1}(n,k)$ and thus (P,[1,n1])(P,[1,n-1]) contains a position with entry $i+1$. To prove (1), Lemma 2.19 assures us that it suffices to prove that (Q,[1,n1])(Q,[1,n-1]) does not contain $i+1$ in any position. Every position in (Q,[1,n1])(Q,[1,n-1]) is either
@@
Since $Q$ is the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$, any position simultaneously in both (Q,[1,n1])(Q,[1,n-1]) and (P,[1,n1])(P,[1,n-1]) cannot contain $i+1$. The position below $P$ contains $i$ since $P$ is a stoplight. Let $R$ be a position between $P$ and $Q$. If the position below $R$ contains $i+1$, then $R$ contains either $i$ or $i+1$. Since $P$ is a stoplight, no position in (P,[1,n1])(P,[1,n-1]) contains $i$, so $R$ cannot contain $i$ (which proves (2)). If $R$ contains $i+1$, then $R$ is not a stoplight because $Q$ is a position in (R,[1,n1])(R,[1,n-1]) with the same value as $R$. So the position below $R$ cannot contain $i+1$, meaning no positions in (Q,[1,n1])(Q,[1,n-1]) contain $i+1$. ∎
@@ 3.2. Proof of homomesy for whirling restricted growth words
Let $g=\mathbf{w}_{\pi}^{-1}(f)$ and suppose $g(i)<g(i+1)$. Let $g(i)=j$. Since $g(i)<g(i+1)$, $g^{-1}(j+1)\not$$\varnothing$. By Propositions 3.4 and 3.5, $f(i)=(\mathbf{w}(g))(i)=(w_{i}(g))(i)\not=1$. ∎
@@ 3.3. Whirling on noncrossing RG-words
Let $f\in\mathcal{F}=\operatorname{RG}_{\mathrm{nc}}(n,k)$ for some $1\leq k\leq n$. If f(n1)=f(n)f(n-1)=f(n), then $(w_{n}(f))(n)=1$.